# Deep Learning-Enabled ISAC-OTFS Pre-equalization Design for Aerial-Terrestrial Networks

### v1

@@ III-A2 MSE Analysis of Demodulated Symbols
MSE of communication symbols is defined as the mean squared error between the DD-domain received communication signal $\mathbf{y}_{C,t}$ and the modulated information symbol vector $\mathbf{s}_{t}$. Mathematically, it is expressed by
@@
where $\beta$ is the power normalization factor used to normalize the power of the received communication signal $\mathbf{y}_{C,t}$ to the power of the modulated information symbol vector $\mathbf{s}_{t}$.
@@ IV Proposed Algorithms
In this section, we present our developed algorithm for solving the optimization problem in (31). We first describe the overall CPBPE framework, which is then followed by a detailed description of the channel prediction module, CSI processing module, and pre-equalization module.
@@ IV-B1 Data Pre-processing
Leveraging the sparsity of OTFS channels, the channel parameters, i.e., delays ${{\tau}_{t}}$, Doppler shifts ${{\nu}_{t}}$ and channel attenuation parameters ${{h}}_{t}$, are adopted as the input of the prediction network instead of the whole channel matrix as considered in [24]. This greatly reduces the dimensionality of the network input, effectively lowering the network overhead. In this way, the communication or sensing channel parameter set at the $t$-th time slot is denoted as
$\mathcal{H}_{t}=\left\{\mathfrak{Re}\left\{{h}_{t}\right\},\mathfrak{Im}\left\{{h}_{t}\right\},{{\tau}_{t}},{{\nu}_{t}}\right\}.$ (33)
@@ IV-B2 Network Structure
Each stack is composed of $B$ blocks. For each block, the input is a channel parameter sequence $\mathcal{H}^{(\zeta)}_{t,\ell}$ of length $\zeta$, containing a portion of the channel features. The output consists of two parts, i.e., one part is a fitted channel parameter sequence of length $\zeta$, representing the features of the channel parameter sequence that the block can fit, and the other part is the predicted channel parameter value ^$\widehat{\mathcal{H}}_{t}$ for the next slot. Note for the first block in the entire model, its input is the original channel parameter sequence $\mathcal{H}^{(\zeta)}_{t}$.
Different blocks are connected based on the doubly residual stacking principle [31]. In other words, each block $\ell$ contains two residual branches, i.e., one branch computes the difference between the fitted channel parameter sequence and the block’s input, outputting the channel parameter sequence that the block failed to fit, which is then passed to the next block for fitting. The other branch sums the predicted channel parameter value ^$\widehat{\mathcal{H}}_{t,\ell}$ of the block $\ell$ with those from other blocks, serving as the channel parameter prediction output for the stack. This process is represented as
@@
^$\widehat{\mathcal{H}}_{t}=\sum_{\ell}\widehat{\mathcal{H}}_{t,\ell}.$ (37)
Inside each block $\ell$, the network is composed of two parts. The first part is a 4-layer fully connected network with a channel parameter sequence $\mathcal{H}^{(\zeta)}_{t,\ell}$ as input. The output of this fully connected network is input into two other fully connected networks. One of these fully connected networks is responsible for fitting the input channel parameter sequence ^$\widehat{\mathcal{H}}^{(\zeta)}_{t,\ell}$ of length $\zeta$. The output result is linearly mapped to output the fitted channel parameter sequence ^$\widehat{\mathcal{H}}^{(\zeta)}_{t,\ell}$. The other fully connected network’s output is linearly mapped to predict the channel parameters ^$\widehat{\mathcal{H}}_{t,\ell}$ for the next time slot. The linear mapping process is expressed as follows:
^$\widehat{\mathcal{H}}_{t,\ell}=\mathbf{V}_{\ell}^{f}\theta_{\ell}^{f}+\mathbf{b}_{\ell}^{f},$ (38)
^$\widehat{\mathcal{H}}^{(\zeta)}_{t,\ell}=\mathbf{V}_{\ell}^{b}\theta_{\ell}^{b}+\mathbf{b}_{\ell}^{b},$ (39)
@@ IV-C CSI Processing Module
TABLE I: DNN network structure
@@
Input layer 1 $(2M\times N,M\times N)$ $[\mathfrak{Re}\{\mathbf{(\hat{H}}_{C,t}^{DD})^{-1}\},\mathfrak{Im}\{\mathbf{(\hat{H}}_{C,t}^{DD})^{-1}\}]$
Input layer 2 $(2M\times N)$ $[\mathfrak{Re}\{\mathbf{s}_{t}\};\mathfrak{Im}\{\mathbf{s}_{t}\}]$
Flatten layer $(2M\times N\times M\times N)$ Input layer 1
Concatenate layer $(2M\times N\times M\times N+2\times M\times N)$ Flatten layer+Input layer 2
Dense layer 1 Tanh $4096$ Concatenate layer
Dense layer 2 Tanh $2048$ Dense layer 1
Dense layer 3 Tanh $1024$ Dense layer 2
Dense layer 4 Tanh $(2M\times N\times M\times N)$ Dense layer 3
Add layer $(2M\times N\times M\times N)$ Flatten layer+Dense layer 4
@@ IV-D1 Data Pre-processing
The inverse matrix of the predicted communication channel is concatenated with the transmitted data symbols with the following processing process: Firstly, the real and imaginary parts of the inverse matrix of the predicted communication channel are separated and reconnected along dimension 1 to form a tensor of size $(2M\times N,M\times N)$. Subsequently, this tensor is flattened into a one-dimensional tensor $\mathbf{t}_{Z}$ of size $(2M\times N\times M\times N)$. Additionally, the real and imaginary parts of the DD-domain modulation symbols $\mathbf{t}_{t}$ are separated and reconnected along dimension 1 to form a one-dimensional tensor $\mathbf{t}_{s}$ of size $(2M\times N)$. Next, $\mathbf{t}_{Z}$ and $\mathbf{t}_{s}$ are concatenated via a concatenate layer to form a one-dimensional tensor $\mathbf{t}_{c}$ of size $(2M\times N\times M\times N+2M\times N)$.
@@ IV-D2 Network Structure
The network consists of multiple dense layers and a residual connection structure. First, $\mathbf{t}_{c}$ is input into the first dense layer through the Tanh activation function. The advantage of using the Tanh activation function lies in its ability to handle negative numbers and maintain a mean value of zero, which is suitable for a pre-equalization matrix that has centralized characteristics. Next, the output of this dense layer is fed into three consecutive dense layers, each with a Tanh activation function. To expedite convergence, we additionally adopt a residual connection method, linking the inverse matrix of the predicted communication channel with the output of the dense layers. More specifically, the output of several consecutive dense layers is connected to the input of the first layer via the residual connection, and the inverse matrix of the predicted communication channel is modified for both communication and sensing performance by directly adding the outputs. Finally, the output is reshaped to size $(2M\times N,M\times N)$ by the reshaping layer. The detailed parameter of the proposed DNN-based pre-equalization design network is listed in TABLE I.
@@ IV-D3 Data Post-processing
The output of the last reshape layer is a matrix with a size of $(2M\times N,M\times N)$, which is denoted as $\bar{\mathbf{P}}_{t}$. The first $MN$ rows of the matrix $\bar{\mathbf{P}}_{t}(1:MN,:)$ correspond to the real part of the pre-equalization matrix before power normalization, and the subsequent $MN$ rows $\bar{\mathbf{P}}_{t}(MN+1:2MN,:)$ correspond to the imaginary part of the pre-equalization matrix before power normalization.
Due to constraints on transmission power in (32), power normalization on the output of the neural network is performed as follows
@@
Then, the real-value outputs are reconnected to the complex-valued pre-equalization matrix as
@@ IV-D4 Implementation Process
Due to the well-established offline training, during the online deployment phase, only one forward propagation is required to complete the prediction of the pre-equalization matrix for the next time slot considering both communication and sensing performance. Therefore, the pre-equalization matrix can be rapidly obtained by inputting the inverse of the channel matrix and the modulated information symbols. After using $\hat{\mathbf{P}}_{t}$ to pre-equalize the transmitted symbols, the sensing performance of the transmitted waveform can be improved, and at the same time, the UE can also directly demodulate the signal without equalization, thus significantly reducing the complexity of the UE.
@@ V Simulation Results
In this section, the numerical results are presented to demonstrate the effectiveness of the proposed scheme. We consider one AP, one UE, which also serves as a sensing target, and $C=2$ scatters. The AP is fixed at coordinates $(0\mathrm{m},0\mathrm{m})$, while the UE and scatterers move around the AP at speeds of up to $25\mathrm{m/s}$. These positions and motion patterns are incorporated into the Wireless Insite [32] software to generate corresponding channels as the dataset, including delays, Doppler shifts, and channel attenuation parameters. The dataset consists of 640 trajectory groups, which are divided into training and testing sets in a ratio of 4:1. Moreover, the size of one OTFS symbol is set as $M=4$ and $N=4$[33, 34], respectively.
To verify the performance of the proposed CPBPE algorithm, we compare our algorithm with the following schemes, i.e.,
@@ V-A Performance of Channel Prediction
As shown in the figure, for the prediction of $\mathfrak{Re}\left(h_{t,p_{C}}\right)$, $\mathfrak{Im}\left(h_{t,p_{C}}\right)$, $\nu_{t,p_{C}}$, and $\tau_{t,p_{C}}$, the NBEATS algorithm always demonstrates the best MAPE performance. The worst prediction performance for the NBESTS algorithm is observed for the delay $\tau_{t,p_{C}}$, but its MAPE is still below 5 %. In addition, for the predictions of $\mathfrak{Re}\left(h_{t,p_{C}}\right)$ and $\mathfrak{Im}\left(h_{t,p_{C}}\right)$, the MAPE is significantly less than 1 %. Overall, this demonstrates that selecting the NBEATS algorithm for OTFS channel parameter prediction is feasible.
Figure 4: Accuracy comparison of OTFS channel parameter prediction.
V-B Effect of Iteration Number
This subsection explores the relationship between the number of iterations during the offline training phase of the pre-equalization design network and the ISAC performance. From Fig. 5, it can be observed that for communication MSE performance, convergence is achieved at approximately 20 iterations. However, for sensing CRLB performance, it takes about 60 iterations to reach convergence. Furthermore, communication MSE initially decreases and then slightly increases during training. This is because, in the early iterations of training, the network primarily focuses on optimizing communication MSE performance. As communication MSE performance stabilizes, the network begins to prioritize optimization of sensing CRLB performance. This indicates that when communication performance is more important in the OTFS-based ISAC system, it may not be necessary to perform a large number of iterations to meet performance requirements. However, if a strict trade-off between communication and sensing performance is desired, more iterations are required. Although the offline training phase takes some iterations, during online deployment, the designed CPBPE network does not require as many iterations and only a single forward propagation is required to complete the design of the pre-equalization matrix considering both communication and sensing performance. The ability of this network to provide rapid outputs helps overcome outdated pre-equalization matrix designs in high-mobility scenarios.

### v2

@@ III-A2 MSE Analysis of Demodulated Symbols
MSE of communication symbols is defined as the mean squared error between the DD-domain received communication signal $\mathbf{y}_{C,t}$ and the modulated symbol vector $\mathbf{s}_{t}$. Mathematically, it is expressed by
@@
where $\beta$ is the power normalization factor used to normalize the power of the received communication signal $\mathbf{y}_{C,t}$ to the power of the modulated symbol vector $\mathbf{s}_{t}$.
@@ IV Proposed Algorithms
In this section, we present our developed algorithm for solving the optimization problem in (27). We first describe the overall CP-PE framework, which is then followed by a detailed description of the channel prediction module, CSI processing module, and pre-equalization module.
@@ IV-B1 Data Pre-processing
Leveraging the sparsity of OTFS channels, the channel parameters, i.e., delays ${\tau}_{t}$}, Doppler shifts ${\nu}_{t}$} and channel attenuation parameters ${{h}}_{t}$, are adopted as the input of the prediction network instead of the whole channel matrix as considered in [26]. This greatly reduces the dimensionality of the network input, effectively lowering the network overhead. In this way, the communication channel parameter set at the $t$-th time slot is denoted as
$\mathcal{H}_{C,t}=\left\{\mathfrak{Re}\left\{{h}_{t}\right\},\mathfrak{Im}\left\{{h}_{t}\right\},{{\tau}_{t}},{{\nu}_{t}}\right\}.$ (29)
Similarly, the sensing channel parameter set can be represented as $\mathcal{H}_{S,t}$. For brevity, the subscripts denoting communication and sensing in $\mathcal{H}_{C,t}$ and $\mathcal{H}_{S,t}$ are omitted in the following.
@@ IV-B2 Network Structure
Each stack is composed of $B$ blocks. For each block, the input is a channel parameter sequence $\mathcal{H}^{(\zeta)}_{t,\ell}$ of length $\zeta$, containing a portion of the channel features. The output consists of two parts, i.e., one part is a fitted channel parameter sequence of length $\zeta$, representing the features of the channel parameter sequence that the block can fit, and the other part is the predicted channel parameter value $\widehat{\mathcal{H}}_{t}$ for the next slot. Note for the first block in the entire model, its input is the original channel parameter sequence $\mathcal{H}^{(\zeta)}_{t}$.
Different blocks are connected based on the doubly residual stacking principle [36]. In other words, each block $\ell$ contains two residual branches, i.e., one branch computes the difference between the fitted channel parameter sequence and the block’s input, outputting the channel parameter sequence that the block failed to fit, which is then passed to the next block for fitting. The other branch sums the predicted channel parameter value $\widehat{\mathcal{H}}_{t,\ell}$ of the block $\ell$ with those from other blocks, serving as the channel parameter prediction output for the stack. This process is represented as
@@
$\widehat{\mathcal{H}}_{t}=\sum_{\ell}\widehat{\mathcal{H}}_{t,\ell}.$ (33)
Inside each block $\ell$, the network is composed of two parts. The first part is a 4-layer fully connected network with a channel parameter sequence $\mathcal{H}^{(\zeta)}_{t,\ell}$ as input. The output of this fully connected network is input into two other fully connected networks. One of these fully connected networks is responsible for fitting the input channel parameter sequence $\widehat{\mathcal{H}}^{(\zeta)}_{t,\ell}$ of length $\zeta$. The output result is linearly mapped to output the fitted channel parameter sequence $\widehat{\mathcal{H}}^{(\zeta)}_{t,\ell}$. The other fully connected network’s output is linearly mapped to predict the channel parameters $\widehat{\mathcal{H}}_{t,\ell}$ for the next time slot. The linear mapping process is expressed as follows:
$\widehat{\mathcal{H}}_{t,\ell}=\mathbf{V}_{\ell}^{f}\theta_{\ell}^{f}+\mathbf{b}_{\ell}^{f},$ (34)
$\widehat{\mathcal{H}}^{(\zeta)}_{t,\ell}=\mathbf{V}_{\ell}^{b}\theta_{\ell}^{b}+\mathbf{b}_{\ell}^{b},$ (35)
@@ IV-C CSI Processing Module
TABLE I: Multi-objective DNN network structure
@@
Input and Pre-processing
Input layer 1 ($\mathbf{H}_{C,t}^{-1}$) $(2M\times N,M\times N)$ $\mathfrak{Re}\{\mathbf{(\hat{H}}_{C,t}^{DD})^{-1}\},\mathfrak{Im}\{\mathbf{(\hat{H}}_{C,t}^{DD})^{-1}\}]$
Input layer 2 ($\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\tau_{p_{S}}}$) $(2M\times N,M\times N)$ $\mathfrak{Re},\mathfrak{Im}]\{\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\tau_{p_{S}}}\}$
Input layer 3 ($\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\nu_{p_{S}}}$) $(2M\times N,M\times N)$ $\mathfrak{Re},\mathfrak{Im}]\{\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\nu_{p_{S}}}\}$
Flatten layers $(2M\times N\times M\times N)$ Each input layer
Communication Branch
Dense layer 1 (comm_dense1) ReLU $128$ Flatten($\mathbf{H}_{C,t}^{-1}$)
Dropout (comm_drop1) $128$ Dense layer 1
Dense layer 2 (comm_dense2) ReLU $256$ Dropout (comm_drop1)
Dropout (comm_drop2) $256$ Dense layer 2
Dense layer 3 (comm_dense3) ReLU $512$ Dropout (comm_drop2)
Dense layer 4 (comm_output) Linear $M\times N\times M\times N$ Dense layer 3
Sensing Branch
Concatenate (sensing_concat) $2\times(2M\times N\times M\times N)$ Flatten($\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\tau_{p_{S}}}$) + Flatten($\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\nu_{p_{S}}}$)
Dense layer 1 (sensing_dense1) ReLU $128$ Concatenate layer
Dropout (sensing_drop1) $128$ Dense layer 1
Dense layer 2 (sensing_dense2) ReLU $256$ Dropout (sensing_drop1)
Dropout (sensing_drop2) $256$ Dense layer 2
Dense layer 3 (sensing_dense3) ReLU $512$ Dropout (sensing_drop2)
Dense layer 4 (sensing_output) Linear $M\times N\times M\times N$ Dense layer 3
Fusion and Output Module
Concatenate (fusion_layer) $2M\times N\times M\times N$ comm_output + sensing_output
Dense layer 1 (fusion_dense1) ReLU $128$ Fusion layer
Dropout (fusion_drop1) $128$ Dense layer 1
Dense layer 2 (fusion_dense2) ReLU $256$ Dropout (fusion_drop1)
Dropout (fusion_drop2) $256$ Dense layer 2
Dense layer 3 (fusion_dense3) ReLU $512$ Dropout (fusion_drop2)
Dense layer 4 (final_output) Linear $2M\times N\times M\times N$ Dense layer 3
Add layer (residual) $2M\times N\times M\times N$ Flatten($\mathbf{H}_{C,t}^{-1}$) + final_output
@@ IV-D1 Data Pre-processing
The input of the proposed multi-objective optimization network consists of the inverse matrix of the predicted communication channel and the sensing derivative matrices $\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\tau_{p_{S}}}$ and $\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\nu_{p_{S}}}$. For each input, the real and imaginary parts are separated and concatenated along dimension 1 to ensure compatibility with real-valued neural network operations. Specifically, the real and imaginary parts of the inverse communication channel matrix are combined to form a tensor of size $(2M\times N,M\times N)$, and each of the derivative matrices $\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\tau_{p_{S}}}$ and $\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\nu_{p_{S}}}$ is processed in the same manner. Each tensor is then flattened into one-dimensional vectors, denoted as $\mathbf{t}_{H}$, $\mathbf{t}_{\tau_{p_{S}}}$, and $\mathbf{t}_{\nu_{p_{S}}}$. These vectors are fed into different branches of the network corresponding to the communication and sensing objectives.
@@ IV-D2 Network Structure
The proposed pre-equalization design network adopts a dual-branch architecture to jointly optimize communication and sensing performance. The communication branch takes the flattened inverse matrix of the predicted communication channel as input and processes it through several dense layers with ReLU activation and dropout for regularization. The sensing branch receives the concatenated flattened vectors of $\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\tau_{p_{S}}}$ and $\dfrac{\partial\mathbf{H}^{p_{S}}}{\partial\nu_{p_{S}}}$, which are then processed through a stack of dense layers. Through this dual-branch architecture, features closely related to pre-equalization are extracted separately from communication and sensing parameters, thereby reducing the processing complexity of subsequent pre-equalization matrix optimization.
Outputs of the two branches are concatenated in a fusion layer, which aggregates features from both domains. The fused features are further processed by dense layers to generate the pre-equalization output. To accelerate convergence and retain model-based priors, a residual connection is introduced between the flattened inverse channel matrix and the output. This enables the network to refine the model-based prior by directly adding learned corrections. Finally, the output is reshaped into a matrix of size $(2M\times N,M\times N)$, corresponding to the real and imaginary parts of the pre-equalization matrix before power normalization. The detailed parameters of the proposed DNN-based pre-equalization design network are listed in TABLE I.
@@ IV-D3 Data Post-processing
The output of the last reshape layer is a matrix with a size of $(2M\times N,M\times N)$, which is denoted as $\bar{\mathbf{P}}_{t}$. The first $MN$ rows $\bar{\mathbf{P}}_{t}(1:MN,:)$ correspond to the real part of the pre-equalization matrix before power normalization, and the subsequent $MN$ rows $\bar{\mathbf{P}}_{t}(MN+1:2MN,:)$ correspond to the imaginary part.
Due to the transmission power constraint in (28), power normalization is applied as
@@
Then, the real-value outputs are reconnected into the complex-valued pre-equalization matrix as
@@ IV-D4 Implementation Process
Due to the well-established offline training, during the online deployment phase, only one forward propagation is required to complete the prediction of the pre-equalization matrix for the next time slot considering both communication and sensing performance. Therefore, the pre-equalization matrix can be rapidly obtained by inputting the inverse of the channel matrix and the modulated symbols. After using $\hat{\mathbf{P}}_{t}$ to pre-equalize the transmitted symbols, the sensing performance of the transmitted waveform can be improved, and at the same time, the UE can also directly demodulate the signal without equalization, thus significantly reducing the complexity of the UE.
@@ V Simulation Results
In this section, the numerical results are presented to demonstrate the effectiveness of the proposed scheme. We consider one AP, one UE, which also serves as a sensing target, and $C=2$ scatters. The AP is fixed at coordinate $(0\mathrm{m},0\mathrm{m})$, while the UE moves around the AP at speeds of up to $95\mathrm{m/s}$. These positions and motion patterns are incorporated into the Wireless Insite [37] software to generate corresponding channels as the dataset, including delays, Doppler shifts, and channel attenuation parameters. The dataset consists of 640 trajectory groups, which are divided into training and testing sets in a ratio of 4:1. Moreover, the size of one OTFS symbol is set as $M=16$ and N=16N=16[38, 39, 40, 41], respectively. The length of the CP is $L=4$. The speed of light is $c=3\times 10^{8}~\mathrm{m/s}$. The carrier frequency is $f_{0}=3~\mathrm{GHz}$, and the subcarrier spacing is $\Delta f=15~\mathrm{kHz}$. Accordingly, the time slot duration is given by $T=66.67~\mu\mathrm{s}$.
To verify the performance of the proposed CP-PE algorithm, we compare our algorithm with the following schemes, i.e.,
@@ V-A Performance of Channel Prediction
As shown in the figure, for the prediction of $\mathfrak{Re}\left(h_{t,p_{C}}\right)$, $\mathfrak{Im}\left(h_{t,p_{C}}\right)$, $\nu_{t,p_{C}}$, and $\tau_{t,p_{C}}$, the NBEATS algorithm always demonstrates the best MAPE performance. The worst prediction performance for the NBESTS algorithm is observed for the delay $\tau_{t,p_{C}}$, while its MAPE is still below 5%. In addition, for the predictions of $\mathfrak{Re}\left(h_{t,p_{C}}\right)$ and $\mathfrak{Im}\left(h_{t,p_{C}}\right)$, the MAPE is significantly less than 1%. Overall, this demonstrates that selecting the NBEATS algorithm for OTFS channel parameter prediction is feasible.
Figure 4: OTFS channel parameter prediction accuracy comparison.
V-B Convergence Behavior
This subsection takes the case where communication and sensing performances are equally weighted (i.e., $\rho_{C}=0.5$, $\rho_{S}=0.5$) as an example and explores the relationship between the number of iterations during the offline training phase of the pre-equalization design network and the ISAC performance. For communication performance, as shown in Fig. 5(a), the MSE gradually converges from a level close to that of the MMSE equalizer to around 0.08. This is because MMSE equalization at the receiver represents the optimal communication performance, while considering sensing performance inevitably leads to a certain degradation in communication MSE. For sensing performance, as shown in Fig. 5(b), $\sqrt{\mathrm{CRLB}_{v}}$ decreases progressively from a value higher than that without pre-equalization to below 0.7 m/s, demonstrating the effectiveness of the proposed pre-equalization design in improving sensing accuracy. Notably, CP-PE can occasionally outperform the case with perfect CSI and sensing parameters, as it optimizes the weighted joint communication–sensing performance. Due to the inherent trade-off between communication and sensing, performance crossover between two algorithms may arise when CRLB is evaluated individually. In both figures, the performance obtained using the outdated pre-equalization matrix is inferior to that based on channel prediction, particularly in terms of communication MSE. This observation indicates that accurate channel information is crucial for the proper operation of the pre-equalization framework.
From the convergence behavior, it can be observed that the training tends to converge after approximately 60 epochs. This convergence mainly corresponds to the optimization of sensing performance, as the communication performance is already close to optimal at the early stage of training. Therefore, when communication performance is prioritized in the OTFS-based ISAC system, a large number of training iterations may not be necessary to meet the performance requirements. However, if a strict trade-off between communication and sensing performance is desired, additional iterations are required to achieve convergence in both objectives. Although the offline training phase takes some iterations, during online deployment, the designed CP-PE network does not require as many iterations and only a single forward propagation is required to complete the design of the pre-equalization matrix considering both communication and sensing performance. The ability of this network to provide rapid outputs helps overcome outdated pre-equalization matrix designs in high-mobility scenarios.